# Verifying LLM Inference to Prevent Model Weight Exfiltration

### v1

@@ 5.2 Scoring Function: IPT Likelihood Score (IPT-LS)
Formally, we compute a verification score for the observed token $t^{\star}$ by integrating a Gaussian density (mean $u$, width $\sigma$) over the token’s probability interval $[a,b]$ in the cumulative distribution $\bar{F}$ of $\vec{p}_{\theta}(x)$. This integral estimates the probability that a perturbed distribution—drawn from $\mathcal{P}_{\theta}(x)$—would produce $t^{\star}$ under the same seed. The log of this mass gives a continuous, differentiable proxy for the fixed-seed posterior probability. The IPT Likelihood Score is defined in Algorithm 3.
Figure 7: Diagram of the IPT Likelihood Score method
5.3 Scoring Function: Gumbel-Max Likelihood Score (GM-LS)
We also introduce the Gumbel-Max Likelihood Score (GM-LS), which is defined for a given noise distribution over token probabilities, which by default we assume to be Gaussian $\mathcal{N}$,
@@ 5.3 Scoring Function: Gumbel-Max Likelihood Score (GM-LS)
Figure 8: Diagram of what scalar the Gumbel-Max Likelihood Score function aims to capture, in order to estimate the likelihood of a particular token being sampled, given a specific noise distribution across token probabilities.
@@ 6.1 Problems with FSSL Estimators and Logit-Rank as a Solution
Our FSSL estimates can be thought of as a “suspiciousness” score. In designing a meaningful score for suspiciousness one prefers that the suspiciousness of a token is a continuous score, from “not-suspicious” to “very suspicious”. However, empirically we find that our proposed scoring function (GM-LS) is not perfectly calibrated. In Qwen-30B experiments, roughly 0.1% of correctly sampled tokens are labeled maximally suspicious, as seen by looking at the middle plot of Figure 9, where we see that the ”rank” of the GM-LS scores is high (in practice, we see that the sampled tokens are tied for last place with a large number of unlikely tokens).
@@
Figure 9: Comparison of Gumbel Likelihood Score (GM-LS) verification and raw logit ranks for sampled tokens. Left: Distribution of GM-LS scores (log-probabilities) for sampled tokens with $\sigma=1.0$. Middle: Distribution of GM-LS ranks, where rank 0 indicates the sampled token has the highest GM-LS score. Ranks are grouped as 0–19 individually and $\geq$20 combined. Right: Distribution of raw logit ranks (before temperature and top-k/top-p filtering), where rank 0 indicates the token with the highest raw logit. All y-axes use log scale to show the full distribution of percentages.
Augmenting GM-LS with a logit rank check yields a strict Pareto improvement over simply thresholding according to the FSSL: for any FSSL threshold $\tau$ and rank cutoff $R$, we introduce a three-class classification
@@
achieves lower or equal FPR at lower information leakage than GM-LS alone. We note that this logit-rank and FSSL score can be combined into a single classifier, e.g. a logistic regression that takes two features: FSSL score and logit-rank. An important advantage to a tiered classification scheme allows to have simpler bounds of extractable information, as there is a trivial upper bound on information expressed by a suspicious (not dangerous) token $I\leq log_{2}(R)\leq log_{2}(|\textrm{vocab size}|)$. As such, given the token-level classifier $C$, the per-token exfiltratable information is:
@@
We sample 500 prompts from LMSYS-Chat-1M, and for each prompt, we generate continuations of max length 500 tokens. We evaluate Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct, Qwen3-1.5-MoE-A.2-7B, and Qwen3-30B-A3B, served via vLLM. [^8] We evaluate on vLLM as this is the most popular open-source inference provider, and thus use GM-LS scores, rather than IPT-LS, as vLLM uses the Gumbel-Max trick for sampling. We use temperature $T=1.$, top-$k=50$, top-$p=0.95$, and a fixed RNG seed for generation. The verifier shares the tokenizer and chat template with the generator. We sweep over multiple $\sigma$ values (from 0.001 to 1.) for GM-LS hyperparameters.
@@
In Section H.1, we demonstrate an ablation of these findings and show a robust improvement when introducing logit rank, where if rank =None then all tokens below an FSSL threshold $\tau$ are classified as dangerous. More information on the GM-LS scores for different tokens are presented in Appendix G.
@@ 8 Acknowledgements
RR was supported in part by the NSF BCS-2218803 grant.
@@ Appendix B Algorithms
B.1 Fixed-Seed Sampling Likelihood Estimators
Algorithm 3 IPT Likelihood Score (from model $\theta$ and context)
@@ B.1 Fixed-Seed Sampling Likelihood Estimators
B.2 Gumbel-Max Likelihood
Algorithm 4 Gumbel-Max Sampled Token Likelihood (GM-STL)
@@ B.3 Naive Methods for Sampled Token Verification
B.4 Rank-Aware GM-LS Scoring
Algorithm 7 Per-Token Risk Classification via GM-LS and Rank
1:token $i$ at position $t$, GM-LS threshold $\tau$, rank cutoff $R$
2:Compute $\mathrm{GM-LS}_{t}(i)$ and $\mathrm{rank}\!\big(\mathrm{logits}_{t}(i)\big)$
3:if $\mathrm{GM-LS}_{t}(i)\geq\tau$ then
@@ D.5 Assumption 5: The randomness used for inference is trusted and non-malleable
If the adversary can manipulate randomness, they can transmit information by selecting seeds/samples while leaving per-step likelihoods within normal ranges. In this regime the verifier’s guarantees degrade sharply: many exfiltration strategies become indistinguishable from benign stochasticity, and a malicious provider could even watermark or backdoor outputs via RNG tampering while remaining white-box undetectable [GKVZ24, CG24].
@@
If an adversary can steer the randomness (e.g., select seeds, bias PRNG state, or grind over retries), the sampler itself becomes a covert channel: the adversary can encode arbitrary payloads by “choosing the random samples.” Closely related phenomena are known in training-time settings, where tampering with algorithmic randomness enables undetectable backdoors that are indistinguishable to any efficient observer [GKVZ24]. Analogously, inference-time coin-fixing enables steganographic signaling without changing logits in a way a verifier could reliably flag (see also pseudo-random watermarking constructions [CG24]).
@@ Appendix E How Specific Implementations may Affect Verification
Another noteworthy consideration is how random samplers update their state, and synchronizing this across the inference and verification servers. For example, in some implementations (like Ollama [Oll]), if you filter the probability distribution with top-p and the token array ends up with only one entry, the inference server will not sample at all, and use a greedy decoding step. This means that the RNG state does not advance for the inference server, but may for the verification server. However, this issue does not arise in other common implementations like vLLM [KLZ+23].
@@ E.1 Speculative Decoding
Speculative decoding [LKM23, CBI+23, ZWHX25, LWZZ25] has emerged as a widely-adopted technique to accelerate LLM inference by using a smaller “draft” model to propose candidate tokens, which are then verified in parallel by the larger “target” model. This approach can achieve substantial speedups (often $1.5{-}2.5\times$) while preserving the exact output distribution of standard autoregressive sampling from the target model.
@@
In speculative decoding introduced by [LKM23, CBI+23], a draft model generates $K$ candidate tokens, which the target model then verifies in parallel. For each candidate, the target model either accepts it, rejects and resamples from an adjusted distribution, or (if all candidates are accepted) samples one additional bonus token. This process maintains exact distributional equivalence to standard autoregressive sampling from the target model alone.
Different speculative decoding algorithms use different methods for accepting tokens from the draft model. For example, in the original work proposing speculative decoding [LKM23] the verifier only accepts a token if it is the exact token that the verifier model would have generated in that position. However, in other methods a draft token is accepted not only if it is the token the verifier would have sampled, but alternatively if it is a sufficiently likely token as well [LWZZ25].
@@ Appendix F Response to Detecting Exfiltration Attempts
This work focuses primarily on the detection of potential model-weight exfiltration attempts rather than on post-detection remediation. In particular, we formalize the task of determining whether a batch of model outputs contains covertly embedded information. Following the literature on steganographic detection [Ker07, KP12], we envision verification pipelines adhering to a general three-step structure:
@@ Appendix G Gumbel-Likelihood Score Statistics
This section presents the empirical distribution of GM-LS scores to provide intuition for how the scoring function behaves on benign traffic and how detection thresholds are calibrated.
@@ G.1 Score Distribution for a Single Token Position
Figure 11 shows the distribution of GM-LS scores across the entire vocabulary for a single token position during generation. For each token $i$ in the vocabulary, we compute $\mathrm{GM-LS}_{t}(i)$ using the Gumbel-Max likelihood estimation procedure described in Algorithm 2.
Figure 11: Distribution of GM-LS scores across vocabulary for a single position. This histogram shows how GM-LS scores are distributed across all possible tokens at a given generation step. The actually-sampled token typically has a score near the peak of this distribution (high likelihood region).
G.2 Distribution of GM-LS Scores for Actually-Generated Tokens
While Figure 11 shows scores for all vocabulary tokens at one position, Figure 12 shows the distribution of GM-LS scores for the tokens that were actually generated across many positions and prompts.
This distribution is crucial for threshold calibration: we generate 500 prompts with up to 500 tokens each under benign conditions (honest model, no adversary), compute $\mathrm{GM-LS}_{t}(y_{t})$ for each actually-generated token $y_{t}$, and collect the resulting scores. The histogram in Figure 12 represents the empirical distribution of GM-LS scores under legitimate operation.
Figure 12: Histogram of GM-LS scores for actually-generated tokens across 500 prompts. Each bin represents how many tokens (out of $\approx$250,000 total) received a given GM-LS score during benign generation. This distribution allows us to select thresholds $\tau$ that achieve target false-positive rates: e.g., setting $\tau$ at the 1st percentile of this distribution yields FPR $\approx 1\%$.
@@ G.3 Threshold Selection and FPR-Information Trade-off
  • Left Y-axis (False Positive Rate): The fraction of benign tokens with $\mathrm{GM-LS}_{t}(y_{t})<\tau$. Computed directly from the cumulative distribution of Figure 12.
  • Right Y-axis (Extractable Information): The worst-case per-token information an adversary can exfiltrate, computed as $\frac{1}{T}\sum_{t}\log_{2}|A_{t}(\tau)|$ where $A_{t}(\tau)$ is the admissible set (tokens with $\mathrm{GM-LS}_{t}(i)\geq\tau$).
@@
Figure 13: Threshold selection trade-off. As the GM-LS threshold $\tau$ increases (moving right), the false positive rate increases (rejecting more benign tokens), while the extractable information decreases (tightening the adversarial capacity bound). This plot allows practitioners to select thresholds that balance security and usability for their deployment.
@@ H.1 Logit-Rank Ablations
Figure 15: Multi-model Pareto frontier with no rank threshold (two-tier baseline). This plot shows the baseline two-tier classifier performance (safe vs. dangerous only, no suspicious category) across all tested models. Without the rank-based stratification, the system can only distinguish between tokens that pass or fail the GM-LS threshold, leading to higher extractable information at comparable FPR levels compared to the three-tier classifier.
Figure 16: Multi-model Pareto frontier with rank threshold = 1. This configuration uses the most aggressive rank cutoff, where only the top-ranked token by raw logits is allowed in the suspicious category. All tokens that fail GM-LS and are not rank-1 are classified as dangerous. This provides the tightest security bounds but may increase the FPR for legitimate low-rank tokens.
Figure 17: Multi-model Pareto frontier with rank threshold = 4. With a rank cutoff of 4, tokens that fail GM-LS but fall within the top 4 by raw logits are classified as suspicious (limited to $\log_{2}(4)=2$ bits of information), while those beyond rank 4 are dangerous. This provides a balance between security and usability.
@@
Figure 20: Qwen1.5-MoE-A2.7B Pareto frontier across rank thresholds. This plot displays the trade-off curves for the Qwen1.5-MoE mixture-of-experts model. Despite having fewer active parameters (2.7B), the MoE architecture shows distinct verification characteristics compared to dense models, providing insights into how model architecture affects the GM-LS verification effectiveness.

### v2

@@ 5.2 Inverse Probability Transform Scoring Function Token-IPT-DiFR
Formally, we compute a verification score for the observed token $t^{\star}$ by integrating a Gaussian density (mean $u$, width $\sigma$) over the token’s probability interval $[a,b]$ in the cumulative distribution $\bar{F}$ of $\vec{p}_{\theta}(x)$. This integral estimates the probability that a perturbed distribution—drawn from $\mathcal{P}_{\theta}(x)$—would produce $t^{\star}$ under the same seed. The log of this mass gives a continuous, differentiable proxy for the fixed-seed posterior probability. The Token-IPT-DiFR is defined in Algorithm 3.
Figure 7: Diagram of the Token-IPT-DiFR method
5.3 Gumbel-Max Scoring Function Token-DiFR
We also introduce the Token-DiFRto estimate FSSL-GM, which is defined for a given noise distribution over token probabilities, which by default we assume to be Gaussian $\mathcal{N}$, also introduced in concurrent work [tokendifr].
@@ 5.3 Gumbel-Max Scoring Function Token-DiFR
Figure 8: Diagram of what scalar the Token-DiFR function aims to capture, in order to estimate the likelihood of a particular token being sampled, given a specific noise distribution across token probabilities.
@@ 6.1 Problems with FSSL Estimators and Logit-Rank as a Solution
Our FSSL estimates can be thought of as a “suspiciousness” score. In designing a meaningful score for suspiciousness one prefers that the suspiciousness of a token is a continuous score, from “not-suspicious” to “very suspicious”. However, empirically we find that our proposed scoring function (FSSL-GM) is not perfectly calibrated. In Qwen-30B experiments, roughly 0.1% of correctly sampled tokens are labeled maximally suspicious, as seen by looking at the middle plot of Figure 9, where we see that the ”rank” of the FSSL-GM scores is high (in practice, we see that the sampled tokens are tied for last place with a large number of unlikely tokens).
@@
Figure 9: Comparison of Token-DiFR(FSSL-GM) verification and raw logit ranks for sampled tokens. Left: Distribution of FSSL-GM scores (log-probabilities) for sampled tokens with $\sigma=1.0$. Middle: Distribution of FSSL-GM ranks, where rank 0 indicates the sampled token has the highest FSSL-GM score. Ranks are grouped as 0–19 individually and $\geq$20 combined; tokens in the $\geq$20 bin are typically tied for last place with many other low-probability tokens (i.e., effective rank $\approx|V|$). Right: Distribution of raw logit ranks (before temperature and top-k/top-p filtering), where rank 0 indicates the token with the highest raw logit. All y-axes use log scale to show the full distribution of percentages.
Augmenting FSSL-GM with a logit rank check yields a strict Pareto improvement over simply thresholding according to the FSSL: for any FSSL threshold $\tau$ and rank cutoff $R$, we introduce a three-class classification
@@
achieves lower or equal FPR at lower information leakage than FSSL-GM alone. We note that this logit-rank and FSSL score can be combined into a single classifier, e.g. a logistic regression that takes two features: FSSL score and logit-rank. An important advantage to a tiered classification scheme allows to have simpler bounds of extractable information, as there is a trivial upper bound on information expressed by a suspicious (not dangerous) token $I\leq log_{2}(R)\leq log_{2}(|\textrm{vocab size}|)$. As such, given the token-level classifier $C$, the per-token exfiltratable information is:
@@
We sample 500 prompts from LMSYS-Chat-1M, and for each prompt, we generate continuations of max length 500 tokens. We evaluate Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct, Qwen3-1.5-MoE-A.2-7B, and Qwen3-30B-A3B, served via vLLM. [^7] We evaluate on vLLM as this is the most popular open-source inference provider, and thus aim toe stimate FSSL-GM scores, rather than FSSL-IPT, as vLLM uses the Gumbel-Max trick for sampling. For all our experiments we use Token-DiFR. We use temperature $T=1.$, top-$k=50$, top-$p=0.95$, and a fixed RNG seed for generation. The verifier shares the tokenizer and chat template with the generator. We sweep over multiple $\sigma$ values (from 0.001 to 1.) for Token-DiFR hyperparameters.
@@
In Section H.1, we demonstrate an ablation of these findings and show a robust improvement when introducing logit rank, where if rank =None then all tokens below an FSSL threshold $\tau$ are classified as dangerous. More information on the FSSL-GM scores for different tokens are presented in Appendix G.
@@ 8 Acknowledgements
Some of this work was conducted while AK, DR, and RR were participating in the ML Alignment and Theory Scholars (MATS) program, whose support we gratefully acknowledge. RR was supported in part by the NSF BCS-2218803 grant, as well as additionally supported by a grant from Coefficient Giving.
@@ Appendix B Algorithms
B.1 Inverse Probability Transform (IPT) Fixed-Seed Sampling Likelihood (FSSL) Estimators
Algorithm 3 Token-IPT-DiFR (from model $\theta$ and context)
@@ B.1 Inverse Probability Transform (IPT) Fixed-Seed Sampling Likelihood (FSSL) Estimators
B.2 Gumbel-Max FSSLEstimators (FSSL-GM)
Algorithm 4 Token-DiFR(used to estimate FSSL-GM)
@@ B.3 Naive Methods for Sampled Token Verification
B.4 Rank-Aware FSSL-GM Scoring
Algorithm 7 Per-Token Risk Classification via FSSL-GM and Rank
1:token $i$ at position $t$, FSSL-GM threshold $\tau$, rank cutoff $R$
2:Compute $\mathrm{FSSL-GM}_{t}(i)$ and $\mathrm{rank}\!\big(\mathrm{logits}_{t}(i)\big)$
3:if $\mathrm{FSSL-GM}_{t}(i)\geq\tau$ then
@@ D.5 Assumption 5: The randomness used for inference is trusted and non-malleable
If the adversary can manipulate randomness, they can transmit information by selecting seeds/samples while leaving per-step likelihoods within normal ranges. In this regime the verifier’s guarantees degrade sharply: many exfiltration strategies become indistinguishable from benign stochasticity, and a malicious provider could even watermark or backdoor outputs via RNG tampering while remaining white-box undetectable [goldwasser2024plantingundetectablebackdoorsmachine, pseudo_random_error_correcting].
@@
If an adversary can steer the randomness (e.g., select seeds, bias PRNG state, or grind over retries), the sampler itself becomes a covert channel: the adversary can encode arbitrary payloads by “choosing the random samples.” Closely related phenomena are known in training-time settings, where tampering with algorithmic randomness enables undetectable backdoors that are indistinguishable to any efficient observer [goldwasser2024plantingundetectablebackdoorsmachine]. Analogously, inference-time coin-fixing enables steganographic signaling without changing logits in a way a verifier could reliably flag (see also pseudo-random watermarking constructions [pseudo_random_error_correcting]).
@@ Appendix E How Specific Implementations may Affect Verification
Another noteworthy consideration is how random samplers update their state, and synchronizing this across the inference and verification servers. For example, in some implementations (like Ollama [ollama_docs]), if you filter the probability distribution with top-p and the token array ends up with only one entry, the inference server will not sample at all, and use a greedy decoding step. This means that the RNG state does not advance for the inference server, but may for the verification server. However, this issue does not arise in other common implementations like vLLM [vllm].
@@ E.1 Speculative Decoding
Speculative decoding [leviathan2023fastinferencetransformersspeculative, chen2023acceleratinglargelanguagemodel, zhang2025learningharmonizedrepresentationsspeculative, li2025eagle3scalinginferenceacceleration] has emerged as a widely-adopted technique to accelerate LLM inference by using a smaller “draft” model to propose candidate tokens, which are then verified in parallel by the larger “target” model. This approach can achieve substantial speedups (often $1.5{-}2.5\times$) while preserving the exact output distribution of standard autoregressive sampling from the target model.
@@
In speculative decoding introduced by [leviathan2023fastinferencetransformersspeculative, chen2023acceleratinglargelanguagemodel], a draft model generates $K$ candidate tokens, which the target model then verifies in parallel. For each candidate, the target model either accepts it, rejects and resamples from an adjusted distribution, or (if all candidates are accepted) samples one additional bonus token. This process maintains exact distributional equivalence to standard autoregressive sampling from the target model alone.
Different speculative decoding algorithms use different methods for accepting tokens from the draft model. For example, in the original work proposing speculative decoding [leviathan2023fastinferencetransformersspeculative] the verifier only accepts a token if it is the exact token that the verifier model would have generated in that position. However, in other methods a draft token is accepted not only if it is the token the verifier would have sampled, but alternatively if it is a sufficiently likely token as well [li2025eagle3scalinginferenceacceleration].
@@ Appendix F Response to Detecting Exfiltration Attempts
This work focuses primarily on the detection of potential model-weight exfiltration attempts rather than on post-detection remediation. In particular, we formalize the task of determining whether a batch of model outputs contains covertly embedded information. Following the literature on steganographic detection [batch_stego_2007, batch_stego_real_world_2012], we envision verification pipelines adhering to a general three-step structure:
@@ Appendix G Gumbel-Likelihood Score Statistics
This section presents the empirical distribution of FSSL-GM scores to provide intuition for how the scoring function behaves on benign traffic and how detection thresholds are calibrated.
@@ G.1 Score Distribution for a Single Token Position
Figure 11 shows the distribution of FSSL-GM scores across the entire vocabulary for a single token position during generation. For each token $i$ in the vocabulary, we compute $\mathrm{FSSL-GM}_{t}(i)$ using the Gumbel-Max likelihood estimation procedure described in Algorithm 2.
Figure 11: Distribution of FSSL-GM scores across vocabulary for a single position. This histogram shows how FSSL-GM scores are distributed across all possible tokens at a given generation step. The actually-sampled token typically has a score near the peak of this distribution (high likelihood region).
G.2 Distribution of FSSL-GM Scores for Actually-Generated Tokens
While Figure 11 shows scores for all vocabulary tokens at one position, Figure 12 shows the distribution of FSSL-GM scores for the tokens that were actually generated across many positions and prompts.
This distribution is crucial for threshold calibration: we generate 500 prompts with up to 500 tokens each under benign conditions (honest model, no adversary), compute $\mathrm{FSSL-GM}_{t}(y_{t})$ for each actually-generated token $y_{t}$, and collect the resulting scores. The histogram in Figure 12 represents the empirical distribution of FSSL-GM scores under legitimate operation.
Figure 12: Histogram of FSSL-GM scores for actually-generated tokens across 500 prompts. Each bin represents how many tokens (out of $\approx$250,000 total) received a given FSSL-GM score during benign generation. This distribution allows us to select thresholds $\tau$ that achieve target false-positive rates: e.g., setting $\tau$ at the 1st percentile of this distribution yields FPR $\approx 1\%$.
@@ G.3 Threshold Selection and FPR-Information Trade-off
  • Left Y-axis (False Positive Rate): The fraction of benign tokens with $\mathrm{FSSL-GM}_{t}(y_{t})<\tau$. Computed directly from the cumulative distribution of Figure 12.
  • Right Y-axis (Extractable Information): The worst-case per-token information an adversary can exfiltrate, computed as $\frac{1}{T}\sum_{t}\log_{2}|A_{t}(\tau)|$ where $A_{t}(\tau)$ is the admissible set (tokens with $\mathrm{FSSL-GM}_{t}(i)\geq\tau$).
@@
Figure 13: Threshold selection trade-off. As the FSSL-GM threshold $\tau$ increases (moving right), the false positive rate increases (rejecting more benign tokens), while the extractable information decreases (tightening the adversarial capacity bound). This plot allows practitioners to select thresholds that balance security and usability for their deployment.
@@ H.1 Logit-Rank Ablations
Figure 15: Multi-model Pareto frontier with no rank threshold (two-tier baseline). This plot shows the baseline two-tier classifier performance (safe vs. dangerous only, no suspicious category) across all tested models. Without the rank-based stratification, the system can only distinguish between tokens that pass or fail the FSSL-GM threshold, leading to higher extractable information at comparable FPR levels compared to the three-tier classifier.
Figure 16: Multi-model Pareto frontier with rank threshold = 1. This configuration uses the most aggressive rank cutoff, where only the top-ranked token by raw logits is allowed in the suspicious category. All tokens that fail FSSL-GM and are not rank-1 are classified as dangerous. This provides the tightest security bounds but may increase the FPR for legitimate low-rank tokens.
Figure 17: Multi-model Pareto frontier with rank threshold = 4. With a rank cutoff of 4, tokens that fail FSSL-GM but fall within the top 4 by raw logits are classified as suspicious (limited to $\log_{2}(4)=2$ bits of information), while those beyond rank 4 are dangerous. This provides a balance between security and usability.
@@
Figure 20: Qwen1.5-MoE-A2.7B Pareto frontier across rank thresholds. This plot displays the trade-off curves for the Qwen1.5-MoE mixture-of-experts model. Despite having fewer active parameters (2.7B), the MoE architecture shows distinct verification characteristics compared to dense models, providing insights into how model architecture affects the FSSL-GM verification effectiveness.